OpenStudy (anonymous):

Rewrite -4(x-4)without parentheses. Explain why each term is positive or negative

OpenStudy (zubhanwc3):

-4x + 16 or -4(x) -4(-4) when multiplied, the outcome is the first equation

OpenStudy (mathstudent55):

Use the distributive property to multiply -4 by x and multiply -4 by 4.

OpenStudy (zubhanwc3):

please close asap.

OpenStudy (mathstudent55):

The distributive property of multiplication over subtraction: \(a(b - c) = ab - ac\) Your problem: \(\color{red}{-4}(\color{green}{x} - \color{orange}{4}) = \color {red}{-4} \cdot \color{green}{x} - \color{red}{(-4)} \cdot \color{orange}{4} = -4x - (-16) = -4x + 16\)